you want to see if there's a way to spend less money on dune buggy rentals. Recall that your current rental car company go beach go charges 29.95 plus 16 cents per mile. nearby are two other companies sunny buggies which charges 72 per day with unlimited free mileage and plava buggies which has no daily fee and charges 45 cents per mile
Answer by rothauserc(4718)

Let x be the number of miles you drive, then the costs of the dune buggy rentals are
:
go Beach go = 0.16x +$29.95
sunny buggies = $72/day
plava buggies = 0.45x
:
suppose your vacation is for three days and you drive 30 miles, then
:
go Beach go = 0.16(30) +$29.95 = $34.75
sunny buggies = $72 * 3 = $216
plava buggies = 0.45 * 30 = $13.50
:
now, suppose you drove 200 miles over the three days
go Beach go = 0.16(200) + $29,95 = $61.95
sunny buggies = $72 * 3 = $216
plava buggies = 0.45 * 200 = $90
:
so the more miles you drive, go Beach go is cheaper BUT if you drive fewer miles then plava buggies is cheaper
: